Example 2: To move charge q from point a to point b requires -30J. Find the voltage drop vh if: (a) q = 6 C, (b) q - -3 C.


1
Expert's answer
2021-09-19T00:23:34-0400

Work Done = W = Voltage Drop * Charge

W = V * Q

Therefore,

"Voltage Drop V = \frac{W}{Q}"


Given W = 30 J


Solution [a]

When Q = 6C

"Therefore, Voltage Drop V = \frac{W}{Q} = \frac{30}{6} = 5V"


Voltage Drop = 5V





Solution [b]

When Q = -3C

"Therefore, Voltage Drop V = \frac{W}{Q} = \frac{30}{-3} = -10 V"


Voltage Drop = -10 V